I need help determining if poisson or exponential would be more appropriate for solving this equation and how you can tell when to apply the exponential. How can you confidently choose a lambda?

Jim and john are fishing. They catch two catfish every five minutes. What is the probability that they wait at least five minutes to catch their first fish?
Answer #1

here when you try to find out probability of first event during an interval ; you should use exponential distribution ;

when you are trying to find out probability of one or more event in an interval you should use Poisson distribution

here \lambda =2/5 =0.4 fish/minute

thereore P( they wait at least five minutes to catch their first fish )=P(T>5)=e-\lambdat =e-2 =0.135335